# Cross-cultural Inspiration Detection and Analysis in Real and LLM-generated Social Media Data

### v1

@@ Abstract
Inspiration is linked to various positive outcomes, such as increased creativity, productivity, and happiness. Although inspiration has great potential, there has been limited effort toward identifying content that is inspiring, as opposed to just engaging or positive. Additionally, most research has concentrated on Western data, with little attention paid to other cultures. This work is the first to study cross-cultural inspiration through machine learning methods. We aim to identify and analyze real and AI-generated cross-cultural inspiring posts. To this end, we compile and make publicly available the InspAIred dataset, which consists of 2,000 real inspiring posts, 2,000 real non-inspiring posts, and 2,000 generated inspiring posts evenly distributed across India and the UK. The real posts are sourced from Reddit, while the generated posts are created using the GPT-4 model. Using this dataset, we conduct extensive computational linguistic analyses to (1) compare inspiring content across cultures, (2) compare AI-generated inspiring posts to real inspiring posts, and (3) determine if detection models can accurately distinguish between inspiring content across cultures and data sources.
@@ 1 Introduction
Despite the compelling motivation, little research has been done on the automatic identification of content that is inspiring, as opposed to simply positive Ignat et al. (2021), and no studies have been conducted on how inspiration varies across different cultures, and whether it can be automatically generated. At the same time, the impressive generative ability of LLMs can create new opportunities for automatically generating inspiring content.
In this work, we aim to address these research gaps by focusing on three main research questions. (RQ1:) How do inspiring posts compare across cultures? (RQ2:) How do AI-generated inspiring posts compare to real inspiring posts across cultures? And (RQ3:) Can detection models effectively differentiate inspiring posts across diverse cultures and data sources?
We summarize our contributions as follows. First, we share a novel dataset comprised of cross-cultural real and generated inspiring and non-inspiring posts, for a total of 2,000 real inspiring posts, 2,000 non-inspiring posts, and 2,000 LLM-generated posts, balanced across India and the UK. Second, we make use of the dataset to conduct extensive computational linguistic analyses to compare inspiring content across cultures and data sources, i.e., India vs. the UK and AI-generated vs. human-written. Finally, we explore the effectiveness of machine learning models for detecting inspiration across diverse data sources and cultures.
@@ 2 Related Work
There have been a limited number of research studies conducted on inspiration, with the majority of them being carried out by the psychology and sociology communities. These studies, such as the ones by Thrash and Elliot (2003, 2004) and Elliot and Thrash (2002), have established the fundamental characteristics of inspiration. Additionally, they have developed a scale to measure the frequency with which people experience inspiration in their daily lives. These studies found that individuals who are inspired tend to be more open to new experiences and show greater absorption in their tasks. They are also more intrinsically motivated, have a strong drive to master their work, and are less competitive.
The work most similar to ours is from Ignat et al. (2021), who are the first to study inspiration through machine learning methods. To facilitate research in this domain, they release a weakly labeled dataset consisting of inspiring and non-inspiring posts collected from Reddit and annotate these posts with their effect on the reader and the emotions they transmit. They also provide a RoBERTa (Liu et al., 2019) classifier fine-tuned on human labels to provide a strong baseline for determining whether a post is inspiring. Finally, they perform extensive data analyses to gain insights into which topics inspire readers and how they influence them. Our work builds on Ignat et al. (2021) by extending it to other cultures, India and the UK, and by collecting AI-generated inspiring posts in order to compare them to human-written posts.
@@
In particular, LLMs excel at creative writing, such as story generation (Yuan et al., 2022), advertising slogan creation (Murakami et al., 2023), and news composition (Yanagi et al., 2020). Tools like Yuan et al. (2022) can help users in their creative pursuits. By generating inspiring content, our work provides a more indirect approach, with the same final goal of helping the user express, transmit, and act on their insight.
@@ 3 The InspAIred Dataset
We use Potato Pei et al. (2022) to create the user interface and store the data. Following the guidelines from Ignat et al. (2021), the user interface contains a definition of inspiration and examples of inspiring and non-inspiring posts, as seen in Figure 2. To find and hire UK annotators, we connect Potato to Prolific.333www.prolific.com We select annotators who have the following qualifications: an approval rate >98%, living in the UK, and a right to vote (as a proxy for age). The annotators independently label 200 posts each, categorizing the posts based on their subjective judgment of whether they found the content inspiring or not. We conduct a similar annotation process with three reliable Indian annotators to label the data from India. To ensure consistency amongst the annotators, they are provided with the same interface as used for the UK annotators and shown in Figure 2.
Finally, we compute the agreement score between the annotators using the Fleiss Kappa measure Fleiss and Cohen (1973). We obtain a score of for UK data and for Indian data, indicating a fair agreement, expected for a subjective task as inspiration detection.
Given that each post is annotated by three annotators, following Ignat et al. (2021), we mark a post as inspiring if at least one annotator labeled it as inspiring.
[…]
| Annotated | Weakly-Labeled | Annotated | Weakly-Labeled | |
[…]
@@ 3.2.1 Prompt Design and Robustness
To ensure the quality of our generated data, we conduct sanity checks to review approximately 200 inspiring posts balanced across the UK and India. The posts are checked for cultural knowledge, factuality, semantic and syntax errors, and style. Based on the feedback, we find that the posts are semantically and syntactically accurate, possess cultural knowledge, and are often more complex than real posts.
@@ 4.1 Stylistic and Structural Features
considers not only word count, but also word complexity. For instance, longer words are more complex than shorter ones. We use the Flesch Reading Ease metric (Flesch, 1948), which counts the number of words per sentence and syllables per word. In Table 3, we find that LLM-generated inspiring posts from the UK and India are less readable than real posts. Additionally, real inspiring posts from India are less readable than those from the UK. Furthermore, when measuring post length, LLM-generated UK posts are longer than real posts and UK posts are shorter than Indian posts.
@@ 4.2 Semantic and Psycholinguistic Features
We use Scattertext Kessler (2017), a tool used to create interactive visualizations of linguistic patterns. After initial pre-processing, we structure the text data for a better representation of information. Next, we use the n-gram representation of the text data to conduct topic extraction through sentence-level clustering. We utilize methods like Non-negative Matrix Factorization (NMF) to decompose the TF-IDF matrix to identify the latent thematic structures within the corpus Ramos et al. (2003).
Results. We analyze the n-gram and topic distributions by various dimensions: inspiring vs. non-inspiring, Indian vs. UK, and real vs. generated. We display the topic distribution across real and generated UK inspiring posts in Figure 3. Further topic and n-gram distributions are displayed in Section A.2.
vs. ✗: Comparing inspiring to non-inspiring real posts, we find that, in the Indian data (Fig. 5, 7), the topic of people is frequently discussed in an inspiring context. Moreover, the most commonly occurring theme in non-inspiring posts is bot, encompassing textual content regarding Reddit rules and moderators.
@@
We use Linguistic Inquiry and Word Count (LIWC), a gold-standard text analysis tool (Pennebaker et al., 2007, 2015), to obtain the words related to human cognitive processes from each post. Specifically, we use the LIWC2015 dictionary, which contains 6,400 words and word stems, each related to a cognitive category. As an example, the word “mother” is assigned the following cognitive categories: female, family, social.
Results. We display the top 24 categories and their LIWC scores, with the most significant differences across dimensions in Table 4. Across all data, we find that male words are more frequent than female, the most common pronoun is I, the most frequent tense is present, and the most common emotion is positive.
vs. ✗: Comparing inspiring to non-inspiring real posts, we find that, in the Indian data, inspiring posts are more likely to be related to socializing, insight, feelings, perception, affection, and contain more positive emotions. Additionally, Indian inspiring posts tend to have more comparisons and use more words related to achievement, health, reward, and work.
@@ 5 Cross-Cultural Inspiration Detection across Real and LLM-generated posts
To answer our last research question – RQ3: Can detection models effectively differentiate inspiring posts across diverse cultures and data sources? – we fine-tune a multi-label classification model to identify if a post is inspiring, represents India or the UK, and whether it is real or LLM-generated.
@@
We use the XLM-RoBERTa base model Conneau et al. (2019), pre-trained on 2.5TB of filtered CommonCrawl Wenzek et al. (2020). We fine-tune the model on our dataset for five epochs with a learning rate of and a batch size of , while monitoring the validation performance and selecting the best model checkpoint based on the F1 metric.
@@ Results.
Since the dataset is evenly distributed across classes, a random baseline results in a 50% accuracy score.
The XLM-RoBERTa model with few-shot setup achieves an accuracy of and F1 score of across all labels, while the default setup achieves an accuracy of and F1 score of . Figure 4 displays the disaggregated results for each label. We find that even with very few training data (600 posts), in the few-shot setup, the XLM-RoBERTa model learns to accurately distinguish inspiring content across cultures (India, UK) and data sources (real and generated). Furthermore, we do not find significant differences in performance across cultures and data sources. Finally, the model can accurately distinguish between real and generated inspiring posts in both few-shot and default training setups.
@@ 6 Conclusion
We hope our work will enable the exploration of various applications to improve creativity and motivation, including storytelling, advertising, and social media, as well as therapy and coaching. Our dataset can be used to test, fine-tune, and analyze other models, and it is publicly available at https://github.com/MichiganNLP/cross_inspiration.
@@
1. A more fine-grained data split
@@
2. Limited number of cross-cultural inspiring posts
@@
3. Relevance of LLM-based data to current times
In our reliance on LLM-based data, it’s imperative to recognize the temporal limitations inherent in its training corpus. While these models offer remarkable capabilities in understanding and generating text, they might not fully capture the current cultural and societal context in their outputs. This limitation can be primarily attributed to the fact that these models are trained on data only up to a specific cutoff date, i.e., GPT-4 only learned from data dated up to September 2021.
@@ Appendix A Appendix
A.1 Fine-tuning Process

### v2

@@ Abstract
Inspiration is linked to various positive outcomes, such as increased creativity, productivity, and happiness. Although inspiration has great potential, there has been limited effort toward identifying content that is inspiring, as opposed to just engaging or positive. Additionally, most research has concentrated on Western data, with little attention paid to other cultures. This work is the first to study cross-cultural inspiration through machine learning methods. We aim to identify and analyze real and AI-generated cross-cultural inspiring posts. To this end, we make publicly available the InspAIred dataset, which consists of 2,000 real inspiring posts, 2,000 real non-inspiring posts, and 2,000 generated inspiring posts evenly distributed across India and the UK. The real posts are sourced from Reddit, while the generated posts are created using the GPT-4 model. Using this dataset, we conduct extensive computational linguistic analyses to (1) compare inspiring content across cultures, (2) compare AI-generated inspiring posts to real inspiring posts, and (3) determine if detection models can accurately distinguish between inspiring content across cultures and data sources. Our dataset can be accessed alongside our classification models at https://github.com/MichiganNLP/cross_inspiration.
@@ 1 Introduction
Despite the compelling motivation, little research has been done on the automatic identification of content that is inspiring, as opposed to simply positive Ignat et al. (2021), and no studies have been conducted on how inspiration varies across cultures, and whether it can be automatically generated. At the same time, LLMs’ impressive generative ability can create new opportunities for automatically generating inspiring content.
In this work, we aim to address these research gaps by focusing on three main research questions. (RQ1:) How do inspiring posts compare across cultures? (RQ2:) How do AI-generated inspiring posts compare to real inspiring posts across cultures? and (RQ3:) Can detection models differentiate inspiring posts across cultures and data sources?
We summarize our contributions as follows. First, we share a novel dataset comprised of cross-cultural real and generated inspiring and non-inspiring posts, for a total of 2,000 real inspiring posts, 2,000 non-inspiring posts, and 2,000 LLM-generated posts, balanced across India and the UK. Second, we make use of the dataset to conduct computational linguistic analyses to compare inspiring content across cultures and data sources, i.e., India vs. the UK and AI-generated vs. human-written. Finally, we explore the effectiveness of machine learning models for detecting inspiration across diverse data sources and cultures.
@@ 2 Related Work
There have been a limited number of research studies conducted on inspiration, with the majority of them being carried out by the psychology and sociology communities. These studies, such as the ones by Thrash and Elliot (2003, 2004) and Elliot and Thrash (2002), have established the fundamental characteristics of inspiration. Additionally, they have developed a scale to measure the frequency with which people experience inspiration in their lives. These studies found that individuals who are inspired tend to be more open to new experiences and show greater absorption in their tasks. They are also more intrinsically motivated, have a strong drive to master their work, and are less competitive.
The work most similar to ours is from Ignat et al. (2021), who are the first to study inspiration through machine learning methods. To facilitate research in this domain, they release a weakly labeled dataset consisting of inspiring and non-inspiring posts collected from Reddit and annotate these posts with their effect on the reader and the emotions they transmit. They also provide a RoBERTa (Liu et al., 2019) classifier fine-tuned on human labels to provide a strong baseline for determining whether a post is inspiring. Finally, they perform data analyses to gain insights into which topics inspire readers and how they influence them. Our work builds on Ignat et al. (2021) by extending it to other cultures, India and the UK, and by collecting AI-generated inspiring posts in order to compare them to human-written posts.
@@
In particular, LLMs excel at creative writing, such as story generation (Yuan et al., 2022), advertising slogan creation (Murakami et al., 2023), and news composition (Yanagi et al., 2020). Tools like Yuan et al. (2022) can help users in their creative pursuits. By generating inspiring content, our work is more indirect but still aims to help the user express and act on their insights.
@@ 3 The InspAIred Dataset
We use Potato Pei et al. (2022) to create the user interface and store the data. Following the guidelines from Ignat et al. (2021), the user interface contains a definition of inspiration and examples of inspiring and non-inspiring posts, as seen in Figure 2. To find and hire UK annotators, we connect Potato to Prolific.333www.prolific.com We select three annotators who have the following qualifications: an approval rate >98%, living in the UK, and a right to vote (as a proxy for age). The annotators independently label 200 posts each, categorizing the posts based on their subjective judgment of whether they found the content inspiring or not. We conduct a similar annotation process with three reliable Indian annotators to label the data from India. To ensure consistency amongst the annotators, they are provided with the same interface as used for the UK annotators and shown in Figure 2.
Finally, we compute the agreement score between the annotators using the Fleiss Kappa measure Fleiss and Cohen (1973). We obtain a score of for UK data and for Indian data, indicating a fair agreement, aligning with Ignat et al. (2021) who obtained a agreement score. Furthermore, by manually examining the low-agreement examples, we find that inspiration is highly subjective and personal, relying on one’s beliefs, interests, or upbringing (e.g., someone not interested in nature will probably not be inspired by David Attenborough). Finally, we make available all the annotations to encourage future work to research the general or personal nature of inspiration based on data and downstream applications.
Given that each post is annotated by three annotators, following Ignat et al. (2021), we mark a post as inspiring if at least one annotator labeled it as inspiring. We do not select the majority label to avoid excluding other annotators’ opinions. Disagreement among annotators in subjective tasks may reflect systematic differences in opinions across groups, not noise (Fleisig et al., 2023).
@@
[…]
| Annotated | Weakly-Labeled | Annotated | Weakly-Labeled | |
[…]
@@ 3.2.1 Prompt Design and Robustness
To ensure the quality of our generated data, we conduct sanity checks to review approximately 200 inspiring posts balanced across the UK and India. The posts are checked for cultural knowledge, factuality, semantic and syntax errors, and style. Based on the feedback, we find that the posts are semantically and syntactically accurate, possess cultural knowledge, and are often more complex than real posts. More information is provided in the Section A.1.
@@
Cost.
We generated 2,000 posts for two cultures, for a total cost of around 50$. (0.03 per 1K input tokens and 0.06 per 1K output tokens).
@@ 4.1 Stylistic and Structural Features
considers not only word count, but also word complexity. For instance, longer words are more complex than shorter ones. We use the Flesch Reading Ease metric (Flesch, 1948), which counts the number of words per sentence and syllables per word. In Table 3, we find that LLM-generated inspiring posts are less readable than real posts. Additionally, real inspiring posts from India are less readable than those from the UK. Furthermore, when measuring post length, LLM-generated UK posts are longer than real posts and UK posts are shorter than Indian posts.
@@ 4.2 Semantic and Psycholinguistic Features
We use Scattertext Kessler (2017) to create interactive visualizations of linguistic patterns. We use the n-gram representation of the text data to conduct topic extraction through sentence-level clustering. Ramos et al. (2003)
Results. We analyze the n-gram and topic distributions by various dimensions: inspiring vs. non-inspiring, Indian vs. UK, and real vs. generated. We display the topic distribution across real and generated UK inspiring posts in Figure 3. Further topic and n-gram distributions are in Section A.3.
vs. ✗: Comparing inspiring to non-inspiring real posts, we find that, in the Indian data (Fig. 5, 7), the topic of people is frequently discussed in an inspiring context. Moreover, the most commonly occurring theme in non-inspiring posts is bot, encompassing content regarding Reddit rules.
@@
We use Linguistic Inquiry and Word Count (LIWC), a popular text analysis tool (Pennebaker et al., 2007, 2015), to obtain the words related to human cognitive processes from each post. Specifically, we use the LIWC2015 dictionary, which contains 6,400 words and word stems, each related to a cognitive category. As an example, the word “mother” is assigned the following cognitive categories: female, family, social. Even though it is less biased than embedding-based models, LIWC has limitations which we discuss in the Limitation section.
Results. We display the top 24 categories and their LIWC scores, with the most significant differences across dimensions in Table 4. We find that male words are more frequent than female, the most common pronoun is I, the most frequent tense is present, and the most common emotion is positive.
vs. ✗: Comparing inspiring to non-inspiring real posts, we find that the Indian inspiring posts are more related to socializing, insight, feelings, perception, affection, and contain more positive emotions. Additionally, Indian inspiring posts tend to have more comparisons and use more words related to achievement, health, reward, and work.
@@ 5 Cross-Cultural Inspiration Detection across Real and LLM-generated posts
To answer our last research question – RQ3: Can detection models effectively differentiate inspiring posts across diverse cultures and data sources? – we fine-tune several classification models to identify if a post is inspiring, represents India or the UK, and whether it is real or LLM-generated.
@@
As a baseline, we use a Random Forest classifier (Ho, 1995) with the default settings from sklearn (Pedregosa et al., 2011) and TF-IDF Ramos et al. (2003) as features.
As a main model, we use the XLM-RoBERTa base model Conneau et al. (2019), pre-trained on 2.5TB of filtered CommonCrawl Wenzek et al. (2020). We fine-tune the model on our dataset for five epochs with a learning rate of and a batch size of . Finally, we also use the pre-trained Llama 2.7b model, from HuggingFace.101010https://huggingface.co/meta-llama/Llama-2-7b We fine-tune the model on our dataset using LoRA Hu et al. (2022) for five epochs with a rank of , learning rate of and a batch size of .111111All the model parameters are shared in the repository.
@@ Results.
Since the dataset is evenly distributed across classes, a random baseline results in a 50% accuracy score. Figure 4 displays the disaggregated results for each label. The Random Forest baseline with TF-IDF scores, trained in a default setup achieves high accuracy scores (), suggesting that important words and topics contribute significantly to inspiration detection across culture and source. This finding is also supported by the topic modeling results in Section 4.2, which shows different topic distributions for each class. Given that Random Forest is an interpretable model, we encourage future work to further analyse the importance of different features for inspiration detection. Finally, we find that even with very few training data (600 posts), in the few-shot setup, the main models, Llama 2.7b and XLM-RoBERTa, learn to accurately distinguish inspiring content across cultures (India, UK) and data sources (real and generated), without significant differences in performance across cultures and data sources.
@@ 6 Conclusion
We hope our work will enable the exploration of various applications to improve creativity and motivation, including storytelling, advertising, social media, therapy, and coaching. Our dataset is available to test, fine-tune, and analyze other models, and it can be accessed alongside our classification models athttps://github.com/MichiganNLP/cross_inspiration.
@@
A more fine-grained data split.
@@
A limited number of cross-cultural inspiring posts.
@@
Data generated with a closed-source model.
We use GPT-4 to generate the inspiring posts, which is not an open-source LLM. We believe it would be valuable for future work to consider generating more data using open-source models like LLaMA or BLOOM. At the same time, we publicly released the data generated with GPT-4 so that others can build on this dataset, and chose this model due to its SOTA performance, worldwide accessibility, and popularity.
Relevance of LLM-based data to current times.
In our reliance on LLM-based data, it’s imperative to recognize the temporal limitations inherent in its training corpus. While these models offer remarkable capabilities in understanding and generating text, they might not fully capture the current cultural and societal context in their outputs. This limitation can be primarily attributed to the fact that these models are trained on data only up to a specific cutoff date, i.e., GPT-4 only learned from data dated up to January 2022.
Limitations of the LIWC-based data analysis.
LIWC is a dictionary where each word is associated with a psycholinguistics category. Therefore, it is less biased than embedding-based models that often contain biased word associations (e.g., man-programmer, woman-homemaker). Based on a simple and scalable word count approach, LIWC is one of the most commonly used data analysis tools. However, it has several limitations. First, it does not consider word sense disambiguation based on context, which can lead to inaccurate results Schwartz et al. (2013). Second, the model was developed and validated for the analysis of long-form writing, which may impact its validity on short social media posts Panger (2016). Third, ongoing research is being conducted to measure the extent to which LIWC demonstrates cross-linguistic and cross-cultural generalizability, given that most studies have been conducted in English Dudău and Sava (2021).
@@ Appendix A Appendix
A.1 Data Quality Assurance
We mention in the paper that the posts are checked for cultural knowledge, factuality, semantic and syntax errors, and style. Specifically, the authors each check 200 real and generated posts from each culture. One author is from India and is, therefore, able to authenticate the cultural and factual aspects of the sampled posts (e.g., Sal Khan and his education contributions, the ISRO, and its scientific progress in India) and check the correctness of the few (<10% of the data) code switch posts between English and Tamil. All the authors possess cultural knowledge about the UK, to recognize the cultural and factual elements of the posts. When unsure, they manually searched Wikipedia for the factuality of the posts (e.g., Sir David Attenborough’s contributions to nature, the British Nicola Adams, the first-ever female boxer to retain an Olympic title).
Finally, all the authors possess proficient English skills to check for semantic and syntax errors in the sampled posts. We find that apart from a few typos (only in real posts), very common in social media posts, the vast majority of the posts (98%) are correct. In terms of style, we observe that LLM-generated sampled posts are more formal and complex than real posts (e.g., there are no typos, and there are more articles and prepositions). Our analysis also confirmed this by automatically measuring the analytic index.
A.2 Fine-tuning Process